# Time-tagging data acquisition system for testing superconducting electronics based on an RFSoC and custom analog frontend R. A. Foster,11footnotetext: Corresponding author.S. Kandeh O. Medeiros

### v1

@@ Abstract
This work demonstrates a cost-effective data acquisition system for electronic circuits that enables more intricate verification techniques than are practical with conventional experimental setups.
@@
The firmware is combined with a thermal-noise-limited analog frontend with over of dynamic range.
Compared to currently available commercial test equipment, this data acquisition system was able to operate a superconducting shift register at a nearly three-times-higher clock frequency ( vs. ) at roughly one seventh the cost.
@@ 1 Introduction
There are many active areas of research to make systems for computation faster and more energy efficient.
For example, one approach is to keep the same system architecture and circuit topologies while focusing on improving the characteristics of the individual transistors by using new materials such as 2D transition-metal-dichalcogenides [1] or carbon nanotubes [2].
Another approach is to use similar system-level architectures, but use different circuits such as those based on superconducting Josephson junctions (JJs) [3, 4]. Other approaches use both different architectures and circuits, such as crossbar arrays of memristors for acceleration of matrix-vector multiplication [5].
With the exception of the memristor work, which used a custom-built data acquisition system for testing, the full circuit in each of these works was operated well below its maximum operating speed, and only a single device or simplified version of the circuit was tested up to a higher speed.
In these specific cases, and in almost every work with novel electronic devices, limitations of test equipment prevent testing of full circuit operation at high speeds.
However, this limitation raises a question of whether the complex circuit can actually work at the speed of the simpler circuit, or if emergent properties of highly-integrated circuits limit their speed.
Standardized interfaces and logic levels of the experimental circuit allow researchers to leverage specialized high-speed test equipment, like bit-error-rate testers.
For example, a low-power superconducting microprocessor based on parametrons was demonstrated at GHz clock speeds [4].
However, even in this case, limitations of the programmability of the test equipment meant that experimenters had to test a single input and single output of their circuit at a time.
Crucially, this meant they could not test the behavior of their circuit while simultaneously toggling the inputs.
@@
In the case of DC-to-SFQ converters, sufficient theoretical understanding of Josephson junctions and the simplicity of the circuit makes it reasonable to assume it works correctly, and more complex testing circuits can progressively be constructed [3]. To address these challenges when testing novel devices, we introduce a cost-effective, high-performance integrated data acquisition system for testing electronic circuits.
@@
Almost all previously demonstrated nanocryotron circuits, including the encoder, were tested with a multichannel oscilloscope and one or more arbitrary waveform generator (AWG) s [8, 12, 13, 14, 15, 16, 17, 11, 18, 19].
This approach works for circuits well for purely combinational inputs and outputs.
However, even with the sophisticated triggering mechanisms of modern digital storage oscilloscopes, testing failure modes of stateful logic can quickly become impossible without saving and post-processing many gigabytes of waveform data.
@@
For example, a commercial-off-the-shelf (COTS)PXI digitizer and AWG were used to test a binary shift register made with superconducting nanocryotrons, reducing the amount of time to perform a bit-error-rate measurement by 50,000 compared to the conventional setup [22].
@@
The Octopux is designed for testing JJ-based circuits, and as a result, its inability to drive high-impedance loads [23] limits its usefulness for testing nanocryotron circuits, where experimental devices can exhibit resistances exceeding when driven out of the superconducting state.
Nanocryotron memory testers have been developed [24, 25], but the overly-specialized firmware limit its applicability to circuits beyond cryogenic memory.
The Quantum Instrumentation Control Kit (QICK) is based on an Radio-Frequency System-on-Chip (RFSoC), and is designed for qubit and frequency-multiplexed detector experiments.
As such, using it for time-domain readout and processing of detectors or complex digital circuits, although technically feasible [26], is not practical.
However, the high sample rate data converters, extensive programmable logic resources, and low noise of the RFSoC platform are still quite advantageous for time-domain experiments.
@@
This work uses the ZCU111 development board with the XCZU28DR, a 1st-gen RFSoC with eight 14-bit radio-frequency digital-to-analog converters (RFDACs) and eight 12-bit radio-frequency analog-to-digital converters (RFADCs).
@@
The data converters are integrated on the SoC and rely on a stripped-down AXI-stream interface for transferring data to and from the programmable logic [27].
@@ 2 Data acquisition system architecture
The RFSoC on the ZCU111 development board contains a multi-core ARM processor capable of running a Linux operating system and is equipped with a high-speed memory interconnect for interfacing with the programmable logic.
@@
(a) Birds-eye-view of experimental setup with liquid helium (LHe) dewar. The custom firmware and software runs on the ZCU111 Radio-Frequency System-on-Chip (RFSoC) development board, which is connected to the dip probe [25] and experimental circuit through a custom analog frontend.
(b) Block diagram of an RFSoC. The processing system contains a multicore ARM processor that runs Linux and communicates with the programmable logic via memory-mapped interfaces using the cache coherent interconnect.
The programmable logic is comprised of memory, lookup tables (LUTs) and flip flops (FFs), dataconverters, and digital signal processing (DSP) modules.
(c) Block diagram of programmable logic as configured for this application.
To communicate with the processing system via the cache coherent interconnect, direct memory access (DMA) is used for bulk transfers of data, whereas status and configuration registers are interfaced with individual reads and writes to short address ranges.
The data source for the radio-frequency digital-to-analog converter (RFDAC) is selectable between a direct digital synthesis (DDS) sinewave generator, a triangle wave generator (TRI), and arbitrary waveform generator (AWG) before it passes through an affine transformation for gain and offset correction.
@@
The samples and corresponding timetags generated by the sample discriminator are saved in a timetagging sample buffer, which can be transferred to main memory via DMA.
The custom data acquisition system is shown in figure 1a, connected to a dip probe [25] housing a superconducting circuit in a liquid helium (LHe) dewar.
@@
High-speed bulk transfers of data are performed with direct memory access (DMA) which uses a dedicated address range separate from the status and configuration registers.
@@ 2.1 Software
The processing system runs a custom Linux image provided by PYNQ that includes the necessary libraries and packages for loading the firmware onto the PL and interfacing with the memory-mapped modules described in the firmware.
@@ 2.2 Field-programmable gate array (FPGA) firmware
Custom field-programmable gate array (FPGA) firmware configures the programmable logic to generate samples for the RFDACs and processes the RFADC data in realtime.
@@
This design choice simplifies timing restrictions for deep pipelines, which can be challenging to design when a full handshake with a ready signal is required [31].
@@ 2.3 Signal generation
The high SFDR was important for characterization of the analog frontend performance to ensure that spurious tones and noise were generated by the analog frontend and not already present in the test signal.
@@ 2.4 Signal acquisition
An additional factor of two in capacity could be gained by using the other on-chip UltraRAM resource.222UltraRAM is a synchronous memory [33], and the sample buffers are implemented as true dual-port asynchronous memories to decouple the read/write bandwidth. However, by wrapping an UltraRAM with a clock-domain-crossing interface on the read port, it could be used in the sample buffer with no changes to the sample buffer logic.
@@
(a) Operating principle of data reduction.
The input data is compared against a high and low threshold to determine when to start and stop saving data.
Input data is passed through a pipeline delay (adjustable by parameter ) to allow samples to be saved before the start event. An output delay (adjustable by ) enables saving data after the stop event.
The grey sections in the output data waveform indicate no valid data is output for that time interval.
(b) Sharing of start/stop events between multiple channels.
The sample discriminator allows start/stop events on one channel to control the output of other channels by changing the event source for the state machine of each channel.
In this example, data from the RFADC is output only when the output of a matched filter is above some threshold.
(e) Diagram of state machine transition rules. The state machine is initialized into the DISABLED state. When in any state other than DISABLED, the output valid signal is high, allowing the output data to be saved. The standard transition order is DISABLEDPRECAPTURECAPTUREDISABLED; the POSTCAPTURE state is only used when a digital trigger is used. The states and criteria for transitions are described in more detail in section 2.4.1.
@@
Eight parallel RFADC channels are processed simultaneously at a clock rate of , corresponding to eight samples per channel per cycle for a sample rate of per channel.
Data signals are represented with thick, black arrows, while the control signals use thin, blue arrows.
@@
Reused submodules, such as the timestamp generator, timer delay, and pipeline delay are highlighted with grey boxes.
@@
These modules provide the start and stop inputs to the finite state machine (FSM) that is also described in figure 2c.
@@ 2.4.1 Zero suppression with realtime sample discriminator
A sample discriminator was implemented in order to remove irrelevant samples and only save voltage waveforms surrounding events of interest.
In order to reconstruct the original waveform, timetags for each event of interest are saved as well.
@@
The sample discriminator uses a hysteretic threshold comparison following the design of [35], however the implementation is different owing to the realtime requirement for this application. Although a threshold-based discriminator may seem limiting, the ability to preprocess signals (e.g. with a filter or more complex signal processing module) before they enter the sample discriminator, combined with the trigger sharing technique illustrated in figure 2b, can enable arbitrarily complex triggering schemes that combine multiple channels.333The current firmware version has a fixed-weight finite-impulse-response filter, so more complicated triggering schemes may require implementation of a custom module.
For example, a matched filter or lightweight neural network444A neural network may be able to handle non-stationary noise better than a matched filter, with the disadvantage of increased implementation complexity. could be implemented, the output of which could be a likelihood estimate of a particular event occurring over the time window of the filter/network.
@@
With the trigger sharing functionality, these likelihood estimates or other higher-order signals could be used to trigger the capture of raw RFADC data, allowing the user to go back and analyze the original data surrounding events of interest after a measurement occurred.
@@
Deletion of unintersting samples by deasserting the output valid signal is illustrated in figure 2a and 2b by the grey-shaded regions in the data out waveform.
The input data is delayed by a runtime configurable amount () to allow the user to save samples before an event of interest occurs.
@@
The asymmetric condition was preferred over a symmetric scheme with a fixed-priority arbiter due the nature of the application: the penalty for accidentally deleting relevant data is considerably higher than the penalty for accidentally saving a little too much data.
As a user, one has several knobs with which they can configure the sample discriminator. First, the user can specify how long data should be saved for before and after an event occurs (, in figure 3). Second, the user can configure which all-below-low and any-above-high events are passed to which state machine, allowing data values on one channel to control capture on another channel ("trigger source" control input in figure 3). Third, one can skip the level-based trigger and supply a digital trigger with an adjustable delay ("is digital" and inputs in figure 3). Finally, if desired, the user can completely bypass the discriminator ("bypass" input in figure 3).
@@ 2.4.2 Event delays: saving data before and after events of interest
The majority of the complexity in the design comes from allowing the user to save samples before the start (any-above-high) event.
@@
The start signal uses a timer delay, which will only delay the final start signal in a burst of multiple start signals that occur in rapid succession.
See figure 2a for an example of two start events in a short window.
@@ 3 Analog frontend characterization
(a) Electrical schematic of analog frontend amplifier chains.
The RFDAC output is buffered by two fully-differential amplifiers to shift the common-mode voltage and convert to single-ended without loading the RFDAC internal biasing circuit, while still providing current down to DC.
The RFADC input is amplified by a low-noise amplifier, converted to a differential signal, and passed through a variable-gain amplifier (VGA) and anti-alias filter.
A high-purity sinusoid is generated with DDS, and the RFADC output is captured with a DMA buffer.
@@
(c) Comparison of SINAD ratio for passive baluns and the active fronted.
The amplifiers improve SINAD by roughly with the floor limited by the thermal noise floor and the noise figure of the TRF37D73 amplifier.
Analog characterization was performed with single-tone loopback tests with the DDS signal generator and a DMA buffer to capture the data from the RFADC.
The goal of this measurement is to quantify the amount of noise and nonlinearity introduced by the receive amplifiers, assuming that the RFDAC and transmit amplifiers can produce a sufficiently clean tone.
The analog frontend (AFE) is compared with a passive breakout board (XM500) with – baluns for single-ended to differential conversion.
Figure 4c shows the SINAD ratio for sinusoidal tones at various frequencies and power levels.
The introduction of the amplifiers in the receive signal path improves SINAD by for small input signals, pushing it close to the thermal noise floor of for bandwidth.
@@
Note that the linearity is limited (as illustrated by the saturation of SINAD for input power levels above ) because the measurement was performed with a fixed setting for the LMH6401 variable-gain amplifier (VGA), so at higher input levels the RFADC input saturated.
The TRF37D73 amplifier has a P1dBin of , so it should be possible to preserve linearity over more than of dynamic range by adjusting the VGA according to the input power.
As shown in figure 4a, the transmit path amplifier contains two stages to step down the common-mode voltage of the RFDAC without loading the on-chip RFDAC biasing network (which degrades its frequency performance).
@@
(b) Electrical diagram of experimental setup. The RFSoCFPGA and analog frontend (AFE) are connected to the sample with an RF probe [25] cooled with liquid helium (see figure 1a for a photo of the experimental setup).
(c) Waveform data comparing a COTS digitizer with a sample rate of (i) to the custom RFSoC digitizer operating at (ii-iii).
@@
Each subplot includes the input stimulus: data and two-phase clock (top, grey) and digitized outputs: shunt1, shunt2 and out (bottom, colored).
@@
When switching from the COTS to the RFSoC digitizer, the sub- rise time of the output signal became visible, and the maximum clock rate the shift register could operate at was increased from to .
(ii) shows operation with a stimulus clock rate of , with the sample discriminator configured with .
@@
Nanocryotrons act as switches to transfer the state between adjacent loops upon the application of a two-phase clock.555
@@
The voltages across the output nTron (out), shift nTron (shunt2), and input nTron (shunt1) were monitored.
For measurements with the COTS system, two external LNA-2500 amplifiers from RF-Bay ( gain total) were used to amplify each output of the shift register (six LNAs required in total) before digitization.
@@
Another phenomenon that becomes observable at higher sample rates is the differentiation of the rising and falling clock edge, which occurs because of excess inductance in series with the nanocryotrons.
This undesirable effect makes it more difficult to determine when the nanocryotron actually switched, and is only exacerbated at higher speeds.
The differentation of clock pulses is most clearly visible in the test with the RFSoC digitizer (middle subplot), where the clock signal uses a rising and falling edge.
@@
Most importantly, however, is the nearly three-fold increase in maximum operating speed of the shift register: with the COTS equipment, the maximum clock rate attainable in experiment was , but when using the data acquisition system presented in this work, the same circuit could operate at , which was the highest frequency at which we observed it operate. At higher clock rates, even for very short clock/input pulses, the shift register would produce incorrect outputs, either switching when it shouldn’t (i.e. 01 errors) or not switching when it should (i.e. 10 errors).
It should be noted that the COTS digitizer system (AWG, digitizer, controller, chassis) as configured cost about seven times more than the ZCU111 RFSoC and custom analog frontend.
The operation of the sample discriminator is shown in the bottom subplot of figure 5c.

### v4

@@ Abstract
This work demonstrates a cost-effective time-domain data acquisition system for electronic circuits that enables more intricate verification techniques than are practical with conventional experimental setups.
@@
The firmware is combined with a thermal-noise-limited analog frontend with of dynamic range.
Compared to currently available commercial test equipment that is seven times more expensive, this data acquisition system was able to operate a superconducting shift register at a nearly three-times-higher clock frequency ( vs. ).
@@ 1 Introduction
Improving the speed and energy efficiency of integrated circuits is an active area of research, inviting approaches that range from device-level improvements [1, 2], to modification at the device- and circuit-level [3, 4], to entirely new processing architectures that leverage unique capabilities of novel devices [5].
With the exception of [5], which used a custom-built data acquisition system for testing, the full circuit in each of these works was operated well below its maximum operating speed, and only a single device or simplified version of the circuit was tested up to a higher speed.
In these specific cases, as is the case for many works involving novel electronic devices, limitations of test equipment can prevent testing of full circuit operation at high speeds.
However, this limitation raises a question of whether the complex circuit can actually work at the speed of the simpler circuit, or if emergent properties of these highly-integrated circuits limit their speed.
Standardized interfaces and logic levels of the experimental circuit allow researchers to leverage specialized high-speed test equipment, like bit-error-rate testers, which helps address this concern.
For example, a low-power superconducting microprocessor based on adiabatic quantum flux parametrons (AQFPs) was demonstrated at gigahertz clock speeds [4].
However, even in this case, limitations of the programmability of the test equipment meant that experimenters could only test a single input and single output of their circuit at a time.
While this provided reassurance that their circuit could operate at high speeds in these specific cases, they unfortunately could not test the behavior of their circuit while simultaneously toggling multiple inputs.
@@
In the case of DC-to-SFQ converters, sufficient theoretical understanding of Josephson junctions and the simplicity of the circuit makes it reasonable to assume it works correctly.
Once these basic self-tests are proven to work, more complex testing circuits can progressively be constructed [3].
To address some of these challenges faced by researchers developing circuits with novel devices that may behave in new or unexpected ways, we introduce a cost-effective, high-performance, integrated data acquisition system for testing electronic circuits, particularly ones with multiple parallel inputs and outputs.
@@
With the exception of a few cases [12, 13, 14, 15], all previously demonstrated nanocryotron circuits, including the encoder, were tested with a multichannel oscilloscope and one or more arbitrary waveform generators (AWGs) [8, 16, 17, 18, 19, 11, 20, 21].
This approach works well for circuits with purely combinational inputs and outputs.
However, even with the sophisticated triggering mechanisms of modern digital storage oscilloscopes, testing failure modes of stateful logic can quickly become impossible without saving and post-processing many tens to hundreds of gigabytes of waveform data.
This is undesirable because more than 90% of the time in an experiment is spent moving data instead of actually measuring the circuit under test.222This 90% figure assumes that acquisition occurs at with a bit depth of and that data transfer is able to saturate a gigabit ethernet connection, which is an optimistic assumption. In practice this may be more like 99% or higher depending on required sampling rate and the ability of the oscilloscope to saturate the network connection.
This reduces the scope of operating margin analyses that can be performed in a reasonable amount of time.
Furthermore, such a large fraction of time moving data instead of measuring the circuit leads to blind time where errors in the circuit may occur without detection.
@@
Such an approach to acquisition allows for faster and more complex measurements, such as bit error rate as a
function of applied magnetic field, temperature, and bias current.
@@
For example, a commercial-off-the-shelf (COTS)PXI digitizer and AWG from Keysight were used to test a binary shift register made with superconducting nanocryotrons, reducing the amount of time to perform a bit-error-rate measurement by 50,000 compared to the conventional setup [13].
@@
The Octopux is designed for testing Josephson junction (JJ)-based circuits, and as a result, its inability to drive high-impedance loads [24] limits its usefulness for testing nanocryotron circuits, where experimental devices can exhibit resistances exceeding when driven out of the superconducting state.
Nanocryotron memory testers have been developed [12, 25], but overly-specialized firmware limits their applicability to circuits beyond cryogenic memory.
Several data acquisition systems based on the Radio-Frequency System-on-Chip (RFSoC) have been demonstrated for frequency-multiplexed readout of quantum and classical superconducting circuits, such as qubits and microwave photon detectors [26, 27, 28].
These solutions leverage the extremely wideband radio-frequency analog-to-digital converters (RFADCs) and radio-frequency digital-to-analog converters (RFDACs) of the RFSoC to control and readout thousands of devices on a single wire.
Although it is possible to bypass the downconversion step in some of these systems to perform time-domain acquisition at baseband, their firmware would need to be modified substantially to capture data at the full sample rate of the RFADCs, which is required for fast nanocryotron devices that can generate pulses shorter than [29].
Another system based on an RFSoC has demonstrated full-rate time-domain readout of superconducting nanowire single photon detectors (SNSPDs) with custom firmware [30].
Due to the lack of readily-available firmware for RFSoCs that fits the requirements for testing nanocryotrons, we chose to develop a custom data acquisition system based on the RFSoC platform due to its low-noise, high sample rate data converters and extensive programmable logic resources.
@@
This work uses the ZCU111 development board with the XCZU28DR, a 1st-gen RFSoC with eight 14-bit RFDACs and eight 12-bit RFADCs.
@@
The data converters are integrated on the SoC and rely on a stripped-down AXI-stream interface for transferring data to and from the field-programmable gate array (FPGA) fabric (programmable logic) [31].
@@ 2 Data acquisition system architecture
The RFSoC on the ZCU111 development board contains two parts: a multi-core ARM processor — the processing system, and a FPGA— the programmable logic.
@@
(a) Birds-eye-view of experimental setup with liquid helium dewar. Custom firmware and software runs on a ZCU111 Radio-Frequency System-on-Chip (RFSoC) development board, which is connected to a dip probe [25] and experimental circuit through a custom analog frontend.
(b) Block diagram of programmable logic as configured for this application.
To communicate with the ARM processor, direct memory access (DMA) is used for bulk transfers of data; status and configuration registers are interfaced with individual reads and writes to short address ranges.
The data source for the radio-frequency digital-to-analog converter (RFDAC) is selectable between a direct digital synthesis (DDS) sinewave generator, a triangle wave generator (TRI), and arbitrary waveform generator (AWG).
@@
The output of the sample discriminator is saved in a buffer that can be transferred to larger off-chip memory via DMA.
The custom data acquisition system is shown in Figure 1a, connected to a dip probe [25] housing a superconducting circuit in a liquid helium dewar.
@@
High-speed bulk transfers of data are performed with DMA, which uses a dedicated address range separate from the status and configuration registers.
@@ 2.1 Software
The processing system runs a custom Linux image provided by PYNQ that includes the necessary libraries and packages for loading the firmware onto the programmable logic and interfacing with the memory-mapped modules described in the firmware.
@@ 2.2 Field-programmable gate array (FPGA) firmware
Custom FPGA firmware configures the programmable logic to generate samples for the RFDACs and processes the RFADC data in realtime.
@@
This design choice simplifies timing restrictions for deep pipelines, which can limit performance and complicate design when a full handshake with a ready signal is required [35].
@@ 2.3 Signal generation
The high SFDR was important for characterization of the analog frontend performance to ensure that any spurious tones and noise measured by the RFADC came from the analog frontend and were not already present in the test signal.
@@ 2.4 Signal acquisition
An additional factor of two in capacity could be gained by using the other on-chip UltraRAM resource.333UltraRAM is a synchronous memory [37], and the sample buffers are implemented as true dual-port asynchronous memories to decouple the read/write bandwidth. However, by wrapping an UltraRAM with a clock-domain-crossing interface on the read port, it could be used in the sample buffer with no changes to the sample buffer logic.
@@
(a) Input data is compared against thresholds, and the current state is tracked to allow for hysteresis.
The output data is delayed to allow samples to be saved before the start event.
The valid signal is pulled low when the discriminator is in the DISABLED state, as indicated by the gray shaded regions on the “data out” waveform.
(b) Sharing of start/stop events between multiple channels is possible by changing the event source for the state machine of each channel.
In this example, we wish to save raw RFADC data only when the output of a matched filter is above some threshold.
(c) Diagram of state machine transition rules. The state machine is initialized into the DISABLED state. When in any state other than DISABLED, the output valid signal is high, allowing the output data to be saved. The states and criteria for transitions are described in more detail in Section 2.4.1.
@@
Eight RFADC channels are processed simultaneously at a clock rate of , corresponding to eight samples per channel per cycle for a sample rate of per channel.
Data signals are represented with thick, black arrows, while the control inputs use thin, blue arrows.
@@
Reused submodules, such as the timestamp generator, timer delay, and pipeline delay are highlighted with gray boxes.
@@
These modules provide the start and stop inputs to the finite state machine (FSM) that is described in Figure 2c.
@@ 2.4.1 Zero suppression with realtime sample discriminator
Similar to the “qualified buffer” of [30], we implemented a sample discriminator based on [39] to remove irrelevant samples and only save voltage waveforms surrounding events of interest.
@@
In order to reconstruct the original waveform, timetags for each event of interest are saved.
Similar to [39] — albeit operating in real-time — our discriminator performs hysteretic comparison against two thresholds, which is useful for ensuring capture of tightly-spaced or small pulses that may occur in a malfunctioning circuit.
Furthermore, our discriminator can share triggers between multiple channels as illustrated in Figure 2b which allows the sample discriminator to trigger capture of a raw signal on a processed copy of that same signal.
The ability to preprocess signals (e.g. with a filter or more complex signal processing module) before they enter the sample discriminator, combined with the trigger sharing technique, can enable arbitrarily complex triggering schemes that combine multiple channels.444The current firmware version has a fixed-weight finite-impulse-response filter, so more complicated triggering schemes would likely require implementation of a custom module.
For example, a matched filter or lightweight neural network555A neural network may be able to handle non-stationary noise better than a matched filter, with the disadvantage of increased implementation complexity. could be implemented, the output of which could be a likelihood estimate of a particular event occurring over the time window of the filter/network.
@@
With the trigger sharing functionality, these likelihood estimates or other higher-order signals could be used to trigger the capture of raw RFADC data, allowing the user to analyze the original data surrounding events of interest after a measurement occurred.
@@
Deletion of uninteresting samples by deassertion of the output valid signal is illustrated in Figure 2a and 2b by the gray-shaded regions in the data out waveform.
The output data is delayed relative to the input data by a runtime configurable amount () to allow the user to save samples before an event of interest occurs.
@@
The asymmetric condition was preferred over a symmetric scheme with a fixed-priority arbiter due the nature of the application: the penalty for accidentally deleting relevant data is considerably higher than the penalty for accidentally saving a few extra samples (e.g. 8) for each event.
As a user, one has several runtime-configurable parameters they can use to customize the operation of the sample discriminator. First, the user can specify how long data should be saved for before and after an event occurs (, in Figure 3). Second, the user can configure which all-below-low and any-above-high events are passed to which state machine, allowing data values on one channel to control capture on another channel ("trigger source" control input in Figure 3). Third, one can skip the level-based trigger and supply a digital trigger with an adjustable delay ("is digital" and inputs in Figure 3). Finally, if desired, the user can completely bypass the discriminator ("bypass" input in Figure 3).
@@ 2.4.2 Event delays: saving data before and after events of interest
The majority of the complexity in the design comes from allowing the user to save a runtime-configurable number of samples before the start (any-above-high) event.
@@
The start signal uses a timer delay, which will only delay the final start signal in a burst of multiple start signals that occur in rapid succession (see Figure 2a for an example of two start events in a short window).
@@ 3 Analog frontend characterization
Analog frontend (AFE) characterization was performed with single-tone loopback tests with the DDS signal generator and a DMA buffer to capture the data from the RFADC.
The goal of this measurement was to quantify the amount of noise and nonlinearity introduced by the receive amplifiers, assuming that the RFDAC and transmit amplifiers can produce a sufficiently clean tone.
The transmit circuit is comprised of two fully-differential amplifiers which convert the RFDAC differential output to single ended without loading its internal biasing circuit.
The receive circuit is comprised of a low-noise amplifier, whose output is converted to a differential signal and passed through a variable-gain amplifier (VGA) and anti-alias filter before being supplied to the RFADC.
The AFE is compared with a passive breakout board (XM500) with – baluns for single-ended to differential conversion.
Figure 4c shows the signal-to-noise and distortion (SINAD) ratio for sinusoidal tones at various frequencies and power levels.
@@
(a) Electrical schematic of analog frontend amplifier chains that perform conversion between single-ended and differential signals.
@@
Testing is performed by injecting a high-purity sinusoid generated with DDS and capturing the RFADC output with a DMA buffer.
@@
The use of the low-noise amplifier in the active frontend improves signal-to-noise and distortion (SINAD) over the passive balun by roughly with the floor limited by the noise figure of the low-noise amplifier.
The introduction of the amplifiers in the receive signal path improves SINAD by for small input signals, pushing it close to the thermal noise floor of for bandwidth as shown in Figure 4c.
@@
Note that the linearity is limited (as illustrated by the saturation of SINAD for input power levels above ) because the measurement was performed with a fixed setting for the LMH6401 VGA, so at higher input levels the RFADC input saturated.
The TRF37D73 amplifier has an input compression power (P1dBin) of , so it should be possible to preserve linearity over more than of dynamic range above the noise floor by adjusting the VGA according to the input power.
As shown in Figure 4a, the transmit path amplifier contains two stages to step down the common-mode voltage of the RFDAC without loading the on-chip RFDAC biasing network (which would degrade its frequency performance).
@@
(b) Electrical diagram of experimental setup. The RFSoCFPGA and AFE are connected to the sample with an RF probe [25] cooled with liquid helium (see Figure 1a for a photo of the experimental setup).
(c) Waveform data comparing a COTS digitizer with a sample rate of (i) to the custom RFSoC digitizer operating at (ii,iii).
@@
Each subplot includes the input stimulus: data and two-phase clock (top, gray) and digitized outputs: shunt1, shunt2 and out (bottom, colored).
@@
When switching from the COTS to the RFSoC digitizer, the sub- rise time of the output signal became visible, and the maximum clock rate the shift register could operate at was increased from to as shown in (iii).
(ii) shows operation of the sample discriminator with .
@@
Nanocryotrons act as switches to transfer the state between adjacent loops upon the application of a two-phase clock.666
@@
The voltages across the output nanocryotron (out), shift nanocryotron (shunt2), and input nanocryotron (shunt1) were monitored.
For measurements with the COTS system from Keysight, two external LNA-2500 amplifiers from RF-Bay ( gain total) were used to amplify each output of the shift register (six LNAs required in total) before digitization.
@@
Another phenomenon that becomes observable
at higher sample rates is the differentiation of the fast clock pulses, which occurs because of excess inductance in series with the nanocryotrons that forces the rapidly-changing current of the clock edge into the shunt resistor. This undesirable effect makes it more difficult to determine when the nanocryotron actually switched, and is only exacerbated at higher clock rates.
The effect is most clearly visible in the test with the RFSoC digitizer 5c(iii), where the clock signal uses a rising and falling edge.
@@
Most importantly, however, is the nearly three-fold increase in maximum operating speed of the shift register: with the COTS equipment, the maximum clock rate attainable in experiment was due to sample rate limitations of the AWG, but when using the data acquisition system presented in this work, the same circuit could operate at , which was the highest frequency at which we observed it operate. At higher clock rates, even for very short clock/input pulses, the shift register would produce incorrect outputs, either switching when it shouldn’t (i.e. 01 errors) or not switching when it should (i.e. 10 errors).
This is not a fundamental limit of nanocryotron devices, which have been demonstrated to operate at speeds exceeding [40] (and individual nanowires have demonstrated speeds of [41]).
The limited operating frequency of the shift register underscores the importance of high-speed test equipment when designing circuits with novel devices: due to the way this particular circuit was designed, it is unable to operate at the maximum speed of the individual devices it is comprised of.
It should be noted that in addition to allowing us to probe the speed limits of this shift register circuit, the COTS digitizer system (AWG, digitizer, controller, chassis) as configured cost about seven times more than the ZCU111 RFSoC and custom analog frontend ($110,000 vs. $16,000).
The operation of the sample discriminator is shown in Figure 5c(iii).